### 04 Mar 2026 v1

# Neuro-Symbolic Financial Reasoning via Deterministic Fact Ledgers and Adversarial Low-Latency Hallucination Detector

Abstract
Standard Retrieval-Augmented Generation (RAG) architectures fail in high-stakes financial domains due to two fundamental limitations: the inherent arithmetic incompetence of Large Language Models (LLMs) and the distributional semantic conflation of dense vector retrieval (e.g., mapping “Net Income” to “Net Sales” due to contextual proximity). In deterministic domains, a 99% accuracy rate yields 0% operational trust. To achieve zero-hallucination financial reasoning, we introduce the Verifiable Numerical Reasoning Agent (VeNRA). VeNRA shifts the RAG paradigm from retrieving probabilistic text to retrieving deterministic variables via a strictly typed Universal Fact Ledger (UFL), mathematically bounded by a novel Double-Lock Grounding algorithm. Recognizing that upstream parsing anomalies inevitably occur, we introduce the VeNRA Sentinel: a 3-billion parameter SLM trained to forensically audit Python execution traces with only one token test budget. To train this model, we avoid traditional generative hallucination datasets in favor of Adversarial Simulation, programmatically sabotaging golden financial records to simulate production-level “Ecological Errors” (e.g., Logic Code Lies and Numeric Neighbor Traps). Finally, to optimize the Sentinel under strict latency budgets, we utilize a single-pass classification paradigm with optional post thinking for debug. We identify the phenomenon of Loss Dilution in Reverse-Chain-of-Thought training and present a novel, OOM-safe Micro-Chunking loss algorithm to stabilize gradients under extreme differential penalization.
1 Introduction
Figure 1: The VeNRA Neuro-Symbolic Paradigm. (Top) Standard RAG suffers from “Stochastic Inaccuracy” due to probabilistic arithmetic and vector conflation. (Bottom) VeNRA parses PDFs into a Universal Fact Ledger (UFL). A Query Navigator retrieves specific variables via a Lexical Gate. An Architect LLM generates a deterministic Python trace ($T$). Crucially, the 33B Sentinel model forensically audits the logic trace before surfacing the answer.
Retrieval-Augmented Generation (RAG) has become the standard for open-domain Question Answering (QA) and is widely regarded as the leading paradigm for grounding AI outputs in source material Dadopoulos et al. (2025). However, its deployment in high-stakes, deterministic domains, specifically financial auditing, remains stalled by a phenomenon we term Stochastic Inaccuracy. In these environments, an accuracy rate of 99% yields 0% operational trust if the 1% error constitutes a sign-convention inversion on a balance sheet or a temporal hallucination in a compliance report. In fact, as PHANTOM demonstrates, financial QA demands exact numerical precision; even minor errors can critically undermine trust Ji et al. (2025).
This friction arises from two fundamental architectural mismatches. First, dense vector retrieval models rely on distributional semantics. Consequently, they frequently map mathematically opposite terms (e.g., “Net Income” vs. “Net Loss”) to proximate vectors due to their identical linguistic contexts, leading to catastrophic retrieval conflation. Basic dense retrieval often fails to fetch the correct evidence in financial documents Dadopoulos et al. (2025), and general-purpose embeddings struggle with finance-specific terms (e.g., “short” vs. “long”) Anderson and Koudas (2024), consistent with our observation of conflating semantically opposite terms. Second, Large Language Models (LLMs) operate as probabilistic next-token predictors, not arithmetic engines. When tasked with financial reasoning, they simulate the syntax of calculation without preserving its mathematical invariants, resulting in plausible-sounding but factually baseless figures Wang et al. (2025). As noted in recent work, LLMs inherently struggle with tasks requiring exact computation Chang et al. (2025), producing outputs that sound plausible but violate arithmetic invariants.
Current methodologies to mitigate these hallucinations rely on two fragile pillars: parametric scaling (utilizing frontier models with higher parameters) and generative evaluation benchmarks (e.g., HaluEval). Both are insufficient for production financial systems. Frontier models introduce prohibitive latency (>2000>2000ms), rendering them unusable for real-time guardrails. In practice, GPT-4 with retrieval often fails on finance questions (81% error/refusal rate) Islam et al. (2023), and feeding longer contexts only yields marginal gains with unacceptable latency Islam et al. (2023). More critically, current evaluation benchmarks rely almost exclusively on Generative Noise, prompting an LLM to invent an error. For example, the HaluEval benchmark generates false answers via an LLM perturbation framework Li et al. (2023), producing semantically obvious, structurally lazy lies (“Strawman Hallucinations”). Real-world financial agents do not fail by inventing fictitious corporate acquisitions; they fail via Ecological Errors: precise mechanical failures such as selecting an adjacent temporal column in a table (Numeric Neighbor Traps) or executing correct Python logic on incorrect variable extractions (Logic Code Lies). Indeed, financial hallucinations often follow specific patterns (numerical miscalculations, temporal shifts, etc.) Tan et al. (2025), which generic benchmarks do not capture.
In this paper, we argue that reliable financial AI cannot be achieved purely through parametric scaling or generative evaluation. Instead, it requires Neuro-Symbolic Cognitive Offloading and Forensic Auditing. We introduce the VeNRA [^1] The code is available in Github: https://github.com/pagand/VeNRA (Verifiable Numerical Reasoning Agent) framework, an end-to-end ecosystem designed to prevent, simulate, and detect financial hallucinations. To mitigate hallucinations, VeNRA replaces unstructured text retrieval with the Universal Fact Ledger (UFL). By strictly extracting text into typed variables backed by Double-Lock Grounding (requiring both character-offset alignment and semantic schema validation), and gating semantic retrieval with deterministic lexical constraints, we relegate the LLM entirely to the role of a deterministic Python code architect. To detect slipped hallucinations, we introduce the VeNRA Sentinel, a locally deployable $3$-B parameter SLM capable of auditing complex mathematical traces in <50<50ms, zero-shot, zero test-time compute. To train the Sentinel, we pioneer Adversarial Simulation, rejecting generative noise in favor of VeNRA-Data: a 10,000-sample dataset constructed by programmatically sabotaging golden financial records to recreate the exact mechanical failures of production RAG systems.
To optimize the Sentinel for zero-latency inference, we adopt a “Label-First” generation paradigm. We identify the phenomenon of Loss Dilution in Reverse-Chain-of-Thought training, where the gradient signal of the verdict token is overwhelmed by the subsequent reasoning tokens. To solve this, and to prevent Vocabulary Out-Of-Memory (OOM) errors caused by massive differential weighting, we present a novel Micro-Chunking Trainer. This algorithm bounds VRAM usage to $\mathcal{O}(c\cdot|V|)$ rather than sequence length, enabling stable training of highly discriminative judges on consumer hardware. In summary, our main contributions are:
  • Neuro-Symbolic Architecture: The Universal Fact Ledger (UFL) with Double-Lock Grounding and Lexical Pre-Filtering, virtually eliminating extraction hallucination and distributional vector conflation.
  • Adversarial Simulation (VeNRA-Data) [^2] https://huggingface.co/datasets/pagand/venra: A novel methodology for generating rigorous hallucination benchmarks via deterministic sabotage (e.g., Logic Code Lies), evaluated via a strict Hybrid-Family splitting protocol to prevent dataset leakage.
  • Forensic Auditing Systems Engineering [^3] https://huggingface.co/spaces/pagand/VeNRA_halDet: The VeNRA Sentinel, a 3B parameter judge achieving low latency verification. We contribute the OOM-safe Micro-Chunking Trainer, mathematically solving the gradient dilution problem in Reverse-CoT architectures.
2 Background and Related Work
Neural RAG has become the standard architecture for knowledge-intensive tasks Lewis et al. (2020). However, its application in high-stakes, deterministic domains such as financial auditing reveals systematic mismatches between probabilistic language modeling and strict numerical reasoning. While existing literature addresses financial QA benchmarks, hallucination detection, and neuro-symbolic execution in isolation, VeNRA integrates these streams to address the specific failure modes of production financial systems.
2.1 Failures of Dense Retrieval in Finance
Classical RAG architectures rely on dense vector retrieval (e.g., DPR, ColBERT) to map queries to evidence Karpukhin et al. (2020); Khattab and Zaharia (2020). While effective for encyclopedic semantic matching, these models suffer from distributional conflation in financial contexts. Embeddings optimized for semantic similarity frequently cluster mathematically opposite terms (e.g., “A owes B” vs. “B owes A”) due to identical linguistic contexts Thakur et al. (2021). Furthermore, recent studies on FinanceBench demonstrate that generic embedding models struggle with the precise temporal and entity alignment required for 10-K filings, where the answer often hinges on distinguishing “2022” from “2021” within dense tabular structures Islam et al. (2023). VeNRA addresses this by shifting from pure semantic retrieval to a hybrid Lexical-Semantic Gate, prioritizing deterministic token overlap for discriminative financial variables.
2.2 Numerical Reasoning and Neuro-Symbolic Agents
Large Language Models (LLMs) act as probabilistic next-token predictors, often failing to maintain arithmetic invariants during text generation Patel et al. (2021). To mitigate this, recent work has pivoted toward neuro-symbolic methods. Chain-of-Thought (CoT) prompting Wei et al. (2022) improves reasoning but remains prone to calculation errors. Program-Aided Language Models (PAL) and Program of Thoughts (PoT) advance this by delegating computation to an external Python interpreter Gao et al. (2023); Chen et al. (2022).
However, benchmarks like FinQA Chen et al. (2021) and TAT-QA Zhu et al. (2021) reveal that even code-generating models struggle with grounding; frequently executing correct logic on hallucinated input variables. VeNRA extends the PAL paradigm by enforcing a stricter separation of concerns: the LLM acts solely as a Code Architect operating over a strictly typed Universal Fact Ledger (UFL), preventing the injection of unsupported scalars into the computation graph.
2.3 Hallucination Benchmarks: Generative vs. Ecological
The evaluation of hallucinations has largely relied on datasets generated via LLM perturbation. Benchmarks like HaluEval Li et al. (2023) and TruthfulQA Lin et al. (2022a) use strong models to rewrite factual sentences into plausible falsehoods. While scalable, this method produces “Semantic Noise”, hallucinations that are often linguistically distinct from the source. In contrast, financial hallucinations are often mechanical: a column shift in a table or a unit swap (e.g., millions to billions). These “Ecological Errors” are underrepresented in generative benchmarks. New benchmarks like FinTextQA and FINDER have been proposed to capture more realistic, expert-level financial queries Dadopoulos et al. (2025); Chen et al. (2024); Choi and Kim (2025), yet they focus primarily on QA accuracy rather than auditing execution traces. VeNRA-Data diverges from prior work by using Adversarial Simulation, deterministic programmatic sabotage of execution traces, rather than generative rewriting, aligning closer to Red Teaming methodologies Ganguli et al. (2022) than traditional QA dataset creation.
2.4 LLM-as-a-Judge and Low-Latency Auditing
Using LLMs as judges to evaluate RAG outputs is a burgeoning field Zheng et al. (2024). However, most “Judge” architectures rely on frontier models (e.g., GPT-4) which incur multi-second latencies, associated cost, privacy concern, inconsistency, prompt sensitivity, bias, etc rendering them unsuitable for real-time guardrails Kim et al. (2023). Furthermore, standard instruction tuning for these judges optimizes for fluent rationales rather than calibrated classification, leading to “sycophancy” Sharma et al. (2023). VeNRA’s Sentinel model contributes to the growing literature on efficient Small Language Models (SLMs) Gunasekar et al. (2023) and single-pass classification. By utilizing a “Label-First” paradigm to tether gradients and a novel Micro-Chunking loss to manage vocabulary memory, we demonstrate that a 3B parameter model can achieve frontier-level auditing performance with single token test-time budget, overcoming the latency-accuracy trade-off inherent in prior LLM-as-a-Judge work.
2.5 The Long-Context Fallacy
Finally, VeNRA challenges the assumption that expanding context windows (e.g., to 128k tokens) solves retrieval challenges. Consistent with the “lost-in-the-middle” phenomenon, LLMs often overlook information buried deep in long contexts Liu et al. (2024), which is critical for footnotes in 10-K filings. VeNRA’s Double-Lock Grounding and Prompt Bookending strategies explicitly mitigate this by structuring the context via the UFL rather than relying on the model’s raw attention span over unstructured text.
3 The VeNRA Architecture: Grounded Neuro-Symbolic RAG
Standard RAG architectures model high-stakes QA as a purely probabilistic sequence generation task. In deterministic domains such as finance, this paradigm induces stochastic inaccuracy: LLMs simulate the syntax of mathematics rather than computing it, and dense vector embeddings frequently conflate mathematically opposite but contextually proximate terms (e.g., “Deferred Revenue” vs. “Realized Revenue”).
To mitigate these fundamental flaws, we introduce VeNRA to shift the RAG paradigm from “Retrieving Text” to “Retrieving Variables”. Rather than forcing a single LLM to simultaneously act as a semantic search engine, an unstructured parser, and an arithmetic calculator, VeNRA strictly decouples these cognitive burdens. Unstructured documents are deterministically parsed into a strongly typed Universal Fact Ledger (UFL), mathematically grounded, and subsequently queried by an LLM constrained strictly to Python code generation.
3.1 Structure-Preserving Ingestion and the Universal Fact Ledger
The Universal Fact Ledger (UFL) serves as the foundational data contract for the VeNRA ecosystem, acting simultaneously as a computational table for arithmetic and a relational graph for semantic traversal. To populate the UFL from unstructured 10-K filings without succumbing to “Context Rot” Hong et al. (2025), the loss of structural lineage and coreference, we employ a deterministic ingestion pipeline.
Trailing-Buffer Semantic Chunking. Standard recursive chunking mechanisms frequently sever cross-boundary coreferences (e.g., a table row referencing “See Note 12”, where the note text resides in the subsequent chunk). To preserve localized context, we developed a newline-preference trailing-buffer algorithm. Let $C_{i}$ be the $i$-th chunk of text, bounded by a character threshold $\tau=3000$. Rather than splitting at an arbitrary character index, the algorithm splits at the most recent newline character. For every generated text block $C_{i}$, a trailing buffer $b_{i-1}$ consisting of the final $k=300$ characters of the preceding block $C_{i-1}$ is injected as a prefix:
$C^{\prime}_{i}=[\text{Previous Context: }b_{i-1}]\oplus C_{i}$ (1)
Crucially, table structures also update this rolling buffer. Because financial table headers carry essential scalar context (e.g., “in millions”), propagating the table’s raw tail into the subsequent prose chunk allows the extraction model to successfully resolve anaphoric scalar references.
Deterministic Table Melting. Generative LLMs exhibit severe unreliability when extracting 2D grid data Zhu et al. (2021). Consequently, VeNRA bypasses LLM extraction for tables entirely, employing deterministic Pandas routines. Raw markdown tables are parsed using a hierarchical disambiguation stack. Indentation markers (e.g., &nbsp;) are computed to generate fully qualified schema paths (e.g., Assets > Current Assets > Inventory), preventing the collapse of distinct sub-categories into ambiguous metrics like “Other.” Furthermore, row-level overrides dynamically intercept global column scale factors; metrics containing substrings such as “per share” enforce a $1.0$ multiplier and a unit classification of USD/Share, ensuring mathematical invariants are preserved.
3.2 Double-Lock Grounding and Post-Hoc Alignment
The primary vulnerability of JSON-constrained LLM extraction is the generation of unsupported numerical values. To solve the “Generator vs. Extractor” dilemma, we force the generative model to provide a grounding_quote ($q$), the verbatim substring from the source chunk ($C$) that justifies the extracted scalar ($v$). Every extracted fact undergoes a rigorous Double-Lock Grounding protocol.
Lock 1: Mechanical Quote Grounding. The algorithm attempts to physically map the quote $q$ to exact character offsets in $C$. Standard normalized sequence matching algorithms (e.g., difflib) over entire text chunks were discarded; we observed that sequence matching against a large denominator (e.g., a 3000-character chunk) asymptotically approaches 0, rendering it empirically useless for isolated numeric verification. Instead, we use a three-tier fallback:
  • Tier 1 (Exact Match): An $\mathcal{O}(n)$ character-exact substring search, followed by a whitespace-normalized pass (collapsing $\setminus s+$ to a single space) to account for LLM-induced spacing artifacts.
  • Tier 2 (Partial Match): The aligner extracts the longest contiguous numeric token $q_{num}$ from $q$ (e.g., “615” from “\$615 million”) and searches for it verbatim, verifying the scalar value independently of surrounding text hallucination.
  • Tier 3 (Fuzzy Sliding-Window): A sliding window $W$ of tokens is passed over $C$. We deliberately avoid Jaccard Similarity, which heavily penalizes the intersection due to the large union of the document window ($|W|\gg|q|$). Instead, we compute a fast token intersection based on Recall: $R=\frac{|q\cap W|}{|q|}$. If $R>0.55$, the window is refined with a localized character-matching algorithm.
If all three tiers fail, the fact is stamped UNALIGNED, confidence is set to $0.0$, and the scalar is restricted from downstream execution.
Lock 2: Semantic Metric Grounding. Mechanical grounding prevents invented numbers, but it does not prevent “Phantom Metrics”, a hallucination where an LLM assigns a verifiably real number to a fabricated metric name. Let $M$ be the set of core tokens in the generated metric name, excluding general financial stop-words (e.g., “net”, “total”, “ratio”). Let $T_{C}$ be the set of tokens in the source chunk. The semantic overlap score is defined as $s=\frac{|M\cap T_{C}|}{|M|}$. If $s<\tau_{sem}$ (where $\tau_{sem}=0.30$), the fact is rejected as a semantic hallucination, successfully preventing the model from assigning orphaned numbers to hallucinatory schemas.
3.3 Schema-Driven Cognitive Offloading
A persistent challenge in applied LLM engineering is “prompt overfitting”, attempting to handle infinite edge cases via increasingly complex system instructions. In finance, edge cases include dynamic formulas (e.g., “floating rate equal to SOFR plus 2.25%”) and covenants (e.g., “leverage ratio not exceeding 3.50”). A naive prompt instructing the LLM to differentiate these leads to downstream mathematical catastrophes when the LLM extracts “2.25” as a fixed interest rate.
VeNRA introduces Schema-Driven Cognitive Offloading. Rather than relying on conditional logic in the prompt, we leverage the LLM’s inherent pre-training in taxonomy by introducing a Pydantic classification enum: fact_type $\in$ {ACTUAL, LIMIT, FORMULA}. Once the model classifies the fact via constrained decoding, deterministic Python post-processing handles the operational logic. For FORMULA types, the system forcefully nullifies the float (num_value = None). This creates a critical safety boundary: when the Code Agent later queries the UFL, it encounters a None type and reads the formula string in the text_nuance metadata, preventing it from executing arithmetic on a partial floating rate. For LIMIT types, the system automatically appends the suffix [Limit] to the metric name, preserving the strict ontological boundary between an actual performance metric and a legal constraint.
3.4 Hybrid Lexical-Semantic Retrieval
Once the Universal Fact Ledger (UFL) is populated, the system must retrieve the precise variables required to satisfy a user query. Pure semantic vector search is fundamentally ill-suited for financial retrieval because dense embedding models (e.g., bge-m3) rely on distributional semantics. Consequently, they frequently group mathematically opposite terms by contextual proximity. A query for “Net Sales” will consistently retrieve chunks discussing “Net Income” with high cosine similarity, leading to cascading arithmetic failures downstream.
To overcome this, VeNRA introduces a Hybrid Lexical-Semantic Retriever, governed by an LLM-based Query Navigator and a deterministic Lexical Gate.
Query Navigation and Schema Mapping. Upon receiving a query, a Navigator SLM evaluates the user intent against a dynamically injected schema summary, which maps canonical entity IDs and high-frequency metric names. The Navigator outputs a bipartite RetrievalPlan containing: (1) a structured UFL filter (e.g., Entity ID, Year, Metric Keywords, and Nuance conditions), and (2) a “Vector Hypothesis”, an idealized text snippet theoretically containing the answer.
Lexical Pre-Filtering (The Semantic Gate). To prevent distributional semantic conflation, all vector retrieval candidates (and fuzzy UFL metric matches) must pass a deterministic Lexical Gate before entering the context window.
Let $Q$ be the multiset of normalized tokens in the query, and $C$ be the multiset of tokens in the candidate chunk. Standard Jaccard Similarity ($J=\frac{|Q\cap C|}{|Q\cup C|}$) is strictly avoided; because financial text chunks are highly verbose ($|C|\gg|Q|$), the massive union denominator forces $J$ toward zero even for perfect semantic matches. Instead, we compute token intersection Recall:
$R_{lex}=\frac{\sum_{t\in\text{set}(Q)}\min(\#(t,Q),\#(t,C))}{|Q|}$ (2)
where #(t,X)\#(t,X) denotes the multiplicity of token $t$ in multiset $X$. If $R_{lex}<\tau_{lex}$ (where $\tau_{lex}=0.30$), the candidate is dropped.
Crucially, two pre-processing steps guarantee the gate’s efficacy:
  1. 1.
    Domain-Specific Stop-Word Filtering: Standard financial modifiers (e.g., “net”, “total”, “per”) are strictly classified as non-discriminative stop-words and removed from $Q$. If “net” were retained, a query for “Net Income” ($Q=\{\text{net},\text{income}\}$) evaluated against a candidate discussing “Net Sales” ($C=\{\text{net},\text{sales}\}$) would yield $R_{lex}=0.50$, falsely passing the threshold. Filtering “net” isolates the discriminative tokens, yielding $R_{lex}=0.0$ and successfully blocking the semantic hallucination.
  2. 2.
[…]
Relational Graph Expansion. Because the UFL acts as a relational graph, retrieval operates along three concurrent expansion pathways:
  1. 1.
    Structured Expansion: The UFL DataFrame is filtered sequentially. If a nuance_focus is generated by the Navigator (e.g., “Restated”), it acts as a boolean AND-gate over the text_nuance column, deterministically isolating adjusted figures.
  2. 2.
    Entity Pivoting: If retrieved UFL rows contain a related_entity_id (e.g., a supplier), the system pivots and queries the vector database for chunks mentioning that specific entity, using a relaxed lexical threshold ($\tau_{lex}=0.20$) to account for sparse entity mentions in prose.
  3. 3.
    Frequency-Based Chunking: To ensure the reasoning agent has access to the qualitative text (the “data hub”) that generated the numbers, the system isolates the most frequent source_chunk_ids from the retrieved UFL rows and forcefully includes their raw text chunks in the context window.
3.5 Program-Aided Execution and Context Assembly
The final architectural phase extends the Program-Aided Language Models (PAL) paradigm Gao et al. (2023). The LLM is stripped of its role as a probabilistic calculator and relegated to the role of a deterministic software architect.
Context Assembly and Priority Scoring. To prevent “Context Rot”, where an LLM loses attention over an expansive context window, the Context Assembler deduplicates all retrieved UFL rows and text chunks. Chunks are ranked via a priority scoring heuristic: a chunk $c$ receives +5+5 points if it is physically linked to a retrieved UFL row (via source_chunk_id), and +1+1 point for every BM25 keyword it contains. The context is strictly truncated to the top $k=5$ chunks and presented as a fused Markdown representation containing both the structured UFL table and the unstructured raw text.
Two-Pass Deterministic Orchestration. VeNRA executes reasoning via a dual-pass neuro-symbolic flow:
  • Pass 1 (The Architect): A reasoning model evaluates the assembled context. It generates a step-by-step logic plan and writes deterministic Python code to compute the answer. The code is executed in an isolated local Python subprocess. If a formula constraint was previously offloaded (Section 3.3) and the UFL num_value is None, the code naturally halts or explicitly requests the missing variable from the text, preventing blind arithmetic.
  • Pass 2 (The Synthesizer): A low-latency SLM ingests the mathematical output from the Python subprocess and formulates the final natural language response, ensuring every stated number is cited directly to a UFL row_id or Chunk ID.
Runtime Fallback Protocol: “Trust Text Over Data”. Despite rigorous ingestion grounding, optical character recognition (OCR), failure in UFL underlying LLM, and PDF parsing anomalies inevitably occur. VeNRA implements a programmatic fault-tolerance mechanism. The Architect agent is explicitly instructed via its system prompt to cross-reference the structured UFL rows against the provided raw text chunks. If the agent detects that the UFL data is malformed (e.g., an unaligned temporal column) or mathematically contradicts the raw text, it programmatically discards the UFL variable. Instead, the Python code is written to extract and compute the values directly from the provided source text string, guaranteeing system resilience.
4 VeNRA-Data: Adversarial Simulation of RAG Failures
To effectively train a low-latency Sentinel model to audit the outputs of the VeNRA architecture, we require a dataset that accurately reflects the failure modes of financial RAG systems. We introduce VeNRA-Data, a heavily curated, 10,000-sample dataset comprising contrastive pairs of grounded and ungrounded financial reasoning traces.
To prevent the classification model from becoming overly dependent on contextual presence, a common flaw where models flag true financial axioms as hallucinations simply because they are omitted from the immediate text, we avoid the binary True/False paradigm. Instead, VeNRA-Data adopts a strict tripartite taxonomy: SUPPORTED (backed by text or valid computation), UNFOUNDED (contradicts evidence, uses ungrounded inputs, or commits logic errors), and GENERAL (axiomatic financial truths independent of the provided context).
4.1 The Flaw in Generative Hallucinations
Current hallucination detection benchmarks (e.g., HaluEval) primarily rely on LLM-generated semantic noise. Dataset creators typically prompt a frontier model to intentionally rewrite a factual sentence into a hallucination. This methodology fundamentally lacks ecological validity. Generative hallucinations tend to be linguistically obvious, structurally lazy, and linguistically distinct from the source text.
Financial RAG systems in production do not fail by inventing linguistically anomalous sentences; they fail via precise mechanical, calculative, and mapping errors. An LLM extracting data from a 10-K is far more likely to select an adjacent temporal column (e.g., $2022$ instead of $2023$) or divide by a hallucinated denominator than it is to invent a fictitious corporate acquisition.
To overcome the limitations of generative noise, we shift the paradigm to Adversarial Simulation. We utilize five high-quality, human-annotated financial QA datasets including FinQA Chen et al. (2021), TAT-QA Zhu et al. (2021), FinanceBench Islam et al. (2023), truthfulQA Lin et al. (2022b), and PHANTOM Ji et al. (2025) as our “Golden” baseline. Rather than asking an LLM to rewrite them, we employ a deterministic Saboteur Engine to systematically mutate specific informational pathways, algorithmically recreating the exact cognitive failures observed in production architectures.
4.2 The Saboteur Engine: Programmatic Injections
The Saboteur Engine consists of modular, deterministic perturbation algorithms. For a given Golden record containing a Query ($Q$), Context ($C$), execution Trace ($T$), and Target Sentence ($S$), the engine attempts to generate an UNFOUNDED “Hard Negative” clone via one of four specific attack vectors.
1. The Logic Code Lie. This vector targets records containing arithmetic traces (e.g., FinQA). It simulates a scenario where the generated Python code is syntactically and mathematically valid, but the input variables hallucinated by the agent are fundamentally incorrect. Let $V_{T}$ be the set of numeric scalars present in the execution trace $T$, and $V_{C}$ be the set of numeric scalars present in the context $C$. The engine isolates a target variable $v\in V_{T}$ (excluding structural indices such as 0 or $1$). It then samples a distractor scalar $d\in\{V_{C}\setminus V_{T}\}$. The trace is mutated via functional substitution: $T^{\prime}=T[v\leftarrow d]$.
Crucially, the Saboteur then executes the hallucinated trace $T^{\prime}$ in a secure sandbox to produce a new resultant scalar $R^{\prime}$. The target sentence is updated to reflect $R^{\prime}$ ($S\leftarrow S^{\prime}$). The resulting record presents a profound challenge to the Judge model: the sentence $S^{\prime}$ perfectly matches the output of the trace $T^{\prime}$, but the trace relies on an input $d$ that misaligns with the semantic text $C$.
2. The Numeric Neighbor Trap. This vector targets tabular extraction failures, simulating the “column/row shift” errors endemic to PDF optical character recognition (OCR) and LLM grid parsing. Let the financial table be represented as a 2D matrix $M$. The engine locates the ground-truth scalar within the matrix at coordinates $(i,j)$ such that $M_{i,j}\in S$. The algorithm deterministically shifts the coordinate space to select a neighboring cell $M^{\prime}$, where $M^{\prime}\in\{M_{i\pm 1,j},M_{i,j\pm 1}\}$.
By substituting $M^{\prime}$ into $S$, the engine simulates two highly prevalent RAG failures: a Temporal Slip (a horizontal column shift, retrieving the wrong fiscal year) or a Metric Slip (a vertical row shift, retrieving an adjacent accounting category).
3. Irrelevancy and Time Warps. This vector simulates staleness and retrieval failure.
  • The Time Warp: If the user query $Q$ contains a specific temporal bounding condition (e.g., the year $2020$), the Saboteur mutates the query $Q\leftarrow Q_{2021}$. The context $C$ and the sentence $S$ remain entirely unchanged. This trains the model to detect scenarios where the provided answer is factually supported by the text, but fails to satisfy the temporal constraints of the user’s prompt.
  • Context Swap (Non-Existence): The engine retains $Q$ and $S$, but replaces the context $C$ entirely with text chunks sampled from a completely disjoint financial document. This creates a “non-existence” hallucination, forcing the Sentinel model to detect when a generated answer is entirely ungrounded by the provided retrieval context.
4. Semantic and Scale Drift. Finally, the engine applies deterministic regex-bounded substitutions to simulate unit and entity mapping failures. Using a predefined domain ontology, scale tokens are inverted (e.g., “millions” $\leftrightarrow$ “billions”), and capitalized Named Entities are swapped with disjoint entities found elsewhere in the text space (e.g., misattributing a subsidiary’s revenue to the parent corporation).
4.3 The Teacher-Auditor Protocol and Noise Injection
While programmatic sabotage deterministically generates hard negatives, it is susceptible to “Accidental Truths”, scenarios where an injected scalar coincidentally matches a valid contextual fact, or where the original “Golden” record itself contains human annotation errors (e.g., “Dirty Gold”).
To guarantee the pristine quality of the training signals, VeNRA-Data employs a multi-tiered Teacher-Auditor Protocol. The Auditor is tasked with independently evaluating the sabotaged tuple $(Q,C,T^{\prime},S^{\prime})$ to verify if it is genuinely unfounded.
Dual-Path Cause-Effect Validation. A critical algorithmic hurdle in automated auditing is the decoupling of cause and effect. In a “Logic Code Lie” (Section 4.2), the Saboteur mutates an input variable ($v\leftarrow d$). However, the Teacher model frequently flags the output scalar (the final computed answer in $S^{\prime}$) as the error span, rather than the injected input $d$. A naive string-matching validation script falsely rejects these valid sabotages because the cause ($d$) does not equal the effect ($S^{\prime}$).
We resolve this via a Dual-Path Validation algorithm. Let $E_{teacher}$ be the error span detected by the Teacher, and $\Theta_{teacher}$ be the internal reasoning trace (the <think> block). A sabotage injection $d$ is validated if it satisfies either of the following conditions:
  1. 1.
    Effect Match: The fuzzy token set ratio between $d$ and $E_{teacher}$ exceeds a threshold $\tau_{fuzz}>60$.
  2. 2.
    Cause Match: The injected scalar $d$ is explicitly present within the Teacher’s causal reasoning block $\Theta_{teacher}$.
Axiomatic Noise Injection. To train the GENERAL class, we utilize TruthfulQA’s finance and economics splits. Initially, these records contain an empty context ($C=\emptyset$). However, training an autoregressive model on this representation induces a spurious correlation: the model learns the structural shortcut $C=\emptyset\implies\texttt{GENERAL}$, bypassing semantic reasoning entirely. To force true logic-awareness, we implement Axiomatic Noise Injection. For every general axiom, we sample a subset of completely irrelevant financial text chunks from a distractor pool ($C_{distractor}$). The model must learn to actively ignore the distractor context and rely on its parametric memory when the trace signals GENERAL_KNOWLEDGE, preventing structural overfitting.
The “Virtual Context” Protocol and HITL Refinement. For highly ambiguous samples where the automated validation yields low confidence, we utilize an AI Proxy Refiner. To minimize context-window bloat and token costs, the Refiner operates under a “Virtual Context” protocol: it is denied access to the raw context $C$. Instead, it relies entirely on the Teacher’s $\Theta_{teacher}$ block as its “eyes.” Guided by a strict meta-prompt, the Refiner surgically corrects nomenclature confusions, sign-convention inversions (e.g., treating a cash outflow as an inflow), and boilerplate evasions, effectively overruling the Teacher model when it exhibits “math-blindness.”. If the result was still ambigious, we tag for human annotation via the specific UI shown in Fig. 2.
Figure 2: The VeNRA Refinement UI. Streamlit UI to faciliate human annotation that are ambiguous from teacher and the AI proxy refiner.
4.4 Contrast Pairs and Hybrid-Family Splitting
The final architectural requirement for VeNRA-Data is the prevention of dataset leakage and the establishment of a causal evaluation framework.
Standard randomized stratification (e.g., an 80/10/10 train/val/test split) over a dataset containing cloned records guarantees severe data leakage. If a Golden record (the Parent) is assigned to the training set, and its sabotaged clone (the Child) is assigned to the validation set, the model evaluates on a context it has already memorized, inflating performance metrics.
To eliminate leakage and maximize the learning signal, we implement Hybrid-Family Splitting. Let a Family $F_{i}$ be defined as a set containing a golden Parent record and its corresponding sabotaged Children: $F_{i}=\{S_{parent},S^{\prime}_{child\_1},\dots\}$. Families are derived programmatically via a lineage heuristic (family_id).
The dataset is partitioned not by individual records, but by entire Families. Families are grouped into four distinct semantic buckets:
  1. 1.
    Sabotage Pairs: Families containing both SUPPORTED and UNFOUNDED siblings.
  2. 2.
    Natural Failures: Families containing organically UNFOUNDED records (e.g., original FinanceBench failures) with no supported sibling.
  3. 3.
    Natural Supported: Families comprising only golden records not selected for sabotage.
  4. 4.
    Axioms: Families representing GENERAL truths.
A stratified split is performed across these buckets, ensuring exactly 10% of the Families from each bucket are allocated to the Validation and Test sets respectively.
This strict topological isolation provides a dual benefit: it mathematically guarantees zero contextual leakage between splits, and it forces the model to learn the delta. By encountering identical contexts and queries during training, where one answer is mathematically sound and the other contains a subtle logic lie, the model’s loss landscape is optimized to detect minute reasoning derivations rather than broad linguistic artifacts. Furthermore, preserving these pairs in the evaluation sets enables the calculation of a strict “Flip Rate” (Sensitivity), measuring the model’s ability to correctly invert its prediction when presented with a minimal adversarial perturbation.
5 VeNRA Sentinel: Forensic Auditing under Latency Constraints
To deploy VeNRA in a production financial environment, the generated answers must be audited in real-time. Frontier LLM-as-a-Judge models introduce prohibitive latency ($\approx 2$-44 seconds) and excessive operational costs. To solve this, we introduce the VeNRA Sentinel, a specialized $3$-B parameter Small Language Model (SLM) based on Qwen-2.5-Coder-3B-Instruct [^4] https://huggingface.co/pagand/venra.
The Sentinel is trained as a “White-Box” hallucination detector. It ingests the User Query ($Q$), the Text Context ($C$), the Python Execution Trace ($T$), and the Target Sentence ($S$), outputting a strict tripartite verdict: SUPPORTED, UNFOUNDED, or GENERAL. To satisfy a strict inference latency budget while maintaining reasoning capabilities, we had to fundamentally alter standard instruction-tuning paradigms.
5.1 The Sytem 1.5 Paradigm and Token Orthogonality
Standard Chain-of-Thought (CoT) prompting dictates that reasoning must precede the final answer (e.g., [Analysis] $\rightarrow$ [Label]). While this allows the model to utilize computational “test time” to arrive at a conclusion, autoregressive generation of a 150-token analysis incurs $\approx 1500$ms of latency on standard consumer hardware, violating our latency constraints.
To satisfy latency constrain, we inspired from SALSA (Single-pass Autoregressive LLM Structured Classification) framework, formatting the training targets as Reverse-CoT:
[…]
In an autoregressive framework, tokens generated after the label cannot mathematically influence the probability distribution of the label itself during a single forward pass. However, training the model to predict the label and immediately explain it acts as a “Gradient Tether.” It forces the hidden states of the Label token to become incredibly dense, requiring it to contain not just the binary classification decision, but the precise latent “seed” of the upcoming forensic explanation. During production inference, we execute a single forward pass, extract the unnormalized logits for the label tokens, and programmatically halt generation, achieving latency verification. If a query is flagged as UNFOUNDED, generation can optionally resume to provide the debug analysis to the end-user.
To extract highly calibrated logits at the very first decoding step, the target tokens must be semantically and mathematically distinct. Naive binary labels (e.g., True / False or Supported / Unfounded) suffer from severe semantic conflation in the model’s pre-training data.
To select optimal classification tokens, we extracted the input embedding matrix $E\in\mathbb{R}^{|V|\times d}$ from the base Qwen model and computed pairwise cosine similarities:
$sim(e_{i},e_{j})=\frac{e_{i}\cdot e_{j}}{\|e_{i}\|\|e_{j}\|}$ (3)
We discovered that standard labels like True / False exhibited a high cosine similarity ($sim\approx 0.68$), meaning their initial vector representations are highly proximate, forcing the attention layers to work harder to separate them. Through programmatic search, we identified a highly orthogonal set of single-token labels: Found, Fake , and General. This set yielded a maximum pairwise cosine similarity of <0.25<0.25, ensuring that the linear classification head operates on easily separable latent vectors.
5.2 The Effects of Loss Dilution
Training the Sentinel model using standard trainers (e.g. SFTTrainer) resulted in catastrophic failure. The model rapidly achieved a low validation loss but exhibited a near-random “Flip Rate” accuracy on the adversarial contrast pairs. The model had learned to be a Sycophant: it minimized global loss by defaulting to the majority class (Found) and generating highly fluent, generic financial analyses.
In a standard autoregressive Cross-Entropy formulation over a sequence of length $N$, the loss is averaged across all active completion tokens:
$\mathcal{L}_{batch}=-\frac{1}{N}\sum_{i=1}^{N}\log P(y_{i}\mid x_{<i})$ (4)
In our Reverse-CoT prompt, the sequence consists of $1$ label token and approximately $150$ analysis tokens. Consequently, the gradient contribution of the critical classification decision is profoundly diluted: $\frac{\partial\mathcal{L}}{\partial\theta_{label}}\propto\frac{1}{151}<0.7\%$. The model is overwhelmingly penalized for grammatical imperfections in its reasoning rather than catastrophic failures in its verdict.
To restore the gradient signal of the verdict, we must apply a large differential weight to the label token. However, implementing a weighted loss via PyTorch’s CrossEntropyLoss(reduction=’none’) on a $3$-B parameter model triggers immediate Out-Of-Memory (OOM) failures on a single GPU (with <50GB of VRAM) [^5] For a micro-batch size $B=2$ and sequence length $S=4096$, calculating the unreduced loss instantiates a tensor of size $B\times S\times|V|]$. At FP32 precision (required for loss stability), this single intermediate tensor consumes $\approx 1.25$ Billion floats ($\approx 5$ GB). Combined with gradients and optimizer states, this vastly exceeds VRAM capacity.. To solve this, we employed a custom Micro-Chunking Trainer. We bypass the model’s internal loss calculation, flattening the shift-logits and shift-labels. We then iteratively slice the sequence into micro-chunks of size $c=512$ entirely within the loss computation loop. For a flat sequence length $T=B\times S$, the weighted loss is accumulated safely:
$\mathcal{L}_{total}=\frac{\sum_{k=0}^{\lceil T/c\rceil}\sum_{i\in C_{k}}w_{i}\mathcal{L}_{CE}(\hat{y}_{i},y_{i})}{\sum_{k=0}^{\lceil T/c\rceil}\sum_{i\in C_{k}}w_{i}}$ (5)
This architectural optimization bounds the VRAM consumption of the loss function to $\mathcal{O}(c\times|V|)$ rather than $\mathcal{O}(B\times S\times|V|)$, maintaining exact mathematical equivalence to the full weighted Cross-Entropy.
5.3 Dynamic Loss Clamping and Differential Weighting
While the Micro-Chunking Trainer enables weighted loss, multiplying the label gradient by $50\times$ introduces severe topological instability (Gradient Seesaw). Imposing a large penalty on a confidently wrong prediction (where $P(\hat{y})\approx 0$) causes the cross-entropy loss to spike asymptotically [^6] During early training, we observed massive gradient norm explosions ($\|g\|>72,000$), causing the model to violently oscillate, updating from predicting $100\%$ Fake in one step to $100\%$ Found in the next.. To stabilize convergence, we introduced Dynamic Loss Clamping, explicitly capping the maximum penalization:
$\mathcal{L}_{weighted}=\min(w_{i}\cdot\mathcal{L}_{CE},\>5.0\cdot w_{max})$ (6)
Initially, we applied the penalty uniformly across all three classes. This induced a secondary failure mode: the model began drastically over-predicting the minority General class. The model utilized the General token as a “safe harbor” to avoid the massive penalization of misclassifying the primary Found/Fake distribution. We resolved this via Asymmetric Differential Weighting: assigning a $50\times$ penalty to Found/Fake to force hard binary logic, but reducing the General penalty to $10\times$, preserving axiomatic knowledge retention without inducing class bleed.
To facilitate training at high adapter ranks ($r=96,128$) under these extreme gradient conditions, we utilized 4-bit NF4 Quantization (QLoRA) paired with Rank-Stabilized LoRA (rsLoRA) Kalajdzievski (2023). By scaling the adapter activations by $\frac{\alpha}{\sqrt{r}}$ rather than $\frac{\alpha}{r}$, rsLoRA prevents gradient vanishing in deeper layers, enabling stable logic-awareness acquisition.
5.4 Prompting Physics and Uncertainty Quantification
Because the base model (Qwen-2.5-Coder-3B-Instruct) possesses a strong inductive bias for code execution, it instinctively attempts to mathematically solve the Python trace provided in the prompt. However, a 3B parameter model lacks the FLOP depth to reliably execute complex 6-digit subtraction in a single forward pass. If it attempts arithmetic simulation, it hallucinates. We counteract this via Algorithmic De-activation in the system prompt [^7] Example prompt: “Do NOT recalculate the math. Assume the arithmetic evaluates to the Claimed Answer. You must verify the EXTRACTION and LOGIC.” to successfully redirects the model’s attention heads from arithmetic simulation toward semantic set-intersection and string-matching.
Causal LLMs suffer from the “Lost in the Middle” phenomenon in large contexts An et al. (2024). To prevent the model from forgetting the verification target after processing 4,000 tokens of 10-K text, we implement Selective Prompt Repetition. We state the Query and Trace at the very top of the prompt, inject the massive context, and strictly repeat the Query and Trace at the bottom boundary immediately preceding the Label: generation token. This aligns with the finding of Leviathan et al. (2025) for non-reasoning tasks.
During inference, raw softmax probabilities are deceptive; a model can output $P_{top1}=0.99$ simply due to softmax temperature, despite high internal confusion Liu et al. (2025). Drawing inspiration from recent production safety frameworks such as HaluGate Park et al. (2025), VeNRA implements a Logit Gap uncertainty threshold: $\Delta=P_{top1}-P_{top2}$. If $\Delta<0.15$, the Sentinel determines the semantic entropy is too high, forcefully overrides the output to Uncertain, and routes the query to a human auditor or a frontier System-2 model.
6 Discussion, Limitations, and Conclusion
6.1 Discussion: The Neuro-Symbolic Efficiency Frontier
The prevailing trajectory in Large Language Model research assumes that hallucination mitigation is a function of parametric scaling. The VeNRA framework demonstrates an alternative, highly efficient paradigm: Neuro-Symbolic Cognitive Offloading. By acknowledging that LLMs are fundamentally probabilistic sequence generators, VeNRA surgically removes their responsibility to perform arithmetic, bounding them instead within strictly typed schemas (the UFL) and deterministic execution environments (Python).
This architectural shift yields a dual benefit. First, it achieves “Zero-Hallucination” arithmetic by construction, provided the initial extraction is grounded. Second, the development of VeNRA-Data proves that the bottleneck in creating reliable “LLM-as-a-Judge” models is not model size, but data ecology. By shifting evaluation benchmarks from linguistically obvious semantic noise to structurally precise Adversarial Simulation (e.g., Logic Code Lies), we empower locally deployable 33B parameter SLMs to achieve forensic verification capabilities previously restricted to frontier models, effectively democratizing high-stakes financial auditing.
6.2 Limitations
Despite its structural rigor, the VeNRA architecture possesses inherent limitations grounded in its reliance on deterministic extraction:
  1. 1.
    Recursive Graph Blindness: The UFL flattens financial reports into a relational table. While this enables rapid SQL-like filtering, it fails to capture recursive second-order relationships (e.g., “Supplier of a Supplier”) inherent in complex supply chains. Currently, the Naive Graph Expansion relies on single-step filtering, potentially missing multi-hop risks.
  2. 2.
    Referential Nuance Loss: Financial footnotes frequently employ referential pointers (e.g., “See Note 12”). While our Trailing-Buffer Chunking captures immediate context, a footnote extracted into the text_nuance column loses its hypertextual integrity. If the referenced note resides in a disjointed chunk not retrieved by the Navigator, the qualitative context is severed, potentially leading to misinterpretation of conditional covenants.
  3. 3.
    The Semantic Schema Bottleneck: The Lexical Pre-Filter effectively prevents vector conflation for standard GAAP metrics. However, highly idiosyncratic corporate taxonomies (e.g., non-GAAP “Adjusted EBITDAR excluding Stock-Based Comp”) may evade the Navigator’s synonym mapping. If the Lexical Gate is too aggressive, it risks filtering out valid but non-standard metric terminologies.
6.3 Conclusion
In high-stakes domains such as finance, a $99\%$ accuracy rate yields $0\%$ operational trust. VeNRA establishes a novel blueprint for grounded Retrieval-Augmented Generation. By pioneering the Universal Fact Ledger (a deterministic variable store), Double-Lock Grounding (a strict extraction firewall), and the VeNRA Sentinel (a latency-optimized forensic judge), we provide a mathematically grounded architecture for reliable financial reasoning. We release the VeNRA-Data adversarial dataset and the OOM-safe Micro-Chunking Trainer to accelerate the development of specialized, logic-aware verification models across the open-source community.
Appendix A Universal Fact Ledger (UFL) 2.0 Schema
The UFL is implemented as a strict Pydantic v2 schema to ensure compatibility with the Python execution engine. It acts as the API contract between the probabilistic extractor and the deterministic solver.
class UFLRow(BaseModel):
    row_id: str                 # md5(entity + metric + period)
    canonical_entity_id: str    # Normalized (e.g., ’ID_AAPL’)
    metric_name: str            # Semantic key (e.g., ’Revenue’)
    # --- Computation & Grounding ---
    num_value: Optional[float]  # None for qualitative facts
    grounding_quote: str        # Verbatim substring for alignment
    unit_normalized: str        # e.g., "USD", "Ratio", "USD/Share"
    scale: float                # Multiplier (e.g., 1e6)
    # --- Context & Filtering ---
    period_end: Optional[str]   # ISO-8601 Date
    period_type: Optional[str]  # ’FY’, ’Q3’, ’TTM’
    doc_section: str            # Breadcrumb path
    source_chunk_id: str        # ChromaDB Foreign Key
    text_nuance: Optional[str]  # Footnotes / formulas
    # --- Alignment Metadata (Post-Hoc Aligner) ---
    char_interval: Optional[Tuple[int, int]]
    alignment_status: Literal["EXACT", "PARTIAL", "FUZZY", "UNALIGNED"]
    confidence_score: float     # 0.95 for Table Melts, 0.0 if Unaligned
Appendix B Adversarial Simulation Protocols (VeNRA-Data)
The Saboteur Engine generates “Hard Negatives” by mutating grounded reasoning traces via four specific attack vectors.
1. Logic Code Lie (Variable Swap)
  • Context: “…revenue was \$100M, while operating expenses were \$80M…”
  • Golden Trace: step_1 = 100 - 80 $\rightarrow$ Target: “Profit was 20.” (SUPPORTED)
  • Sabotaged Trace: step_1 = 150 - 80 $\rightarrow$ Target: “Profit was 70.” (UNFOUNDED)
  • Task: Judge must detect that 150 does not match the text context 100.
2. Numeric Neighbor Trap (Table Shift)
  • Context: A matrix where $M_{2022}=45.5$ and $M_{2023}=52.1$.
  • Query: “What was the metric in 2022?”
  • Sabotaged Sentence: “The metric in 2022 was 52.1.” (UNFOUNDED: Temporal slip).
3. Time Warp (Irrelevancy)
  • Golden Query: “What was Apple’s revenue in 2021?”
  • Sabotaged Query: “What was Apple’s revenue in 2020?”
  • Note: Context and Answer remain entirely unchanged. The model must recognize that the correct evidence answers the wrong temporal question.
The Teacher-Auditor Validation Protocol To prevent “Accidental Truths” (where a sabotage coincidentally matches reality), we employ a Dual-Path Validation check using a GPT-4o Auditor:
  • Effect Match: Does the Auditor’s detected error span fuzzy-match our injected value?
  • Cause Match: Does the injected value appear explicitly in the Auditor’s internal reasoning trace ($\Theta_{teacher}$)?
Only samples satisfying one of these conditions are accepted into the dataset.
Appendix C Micro-Chunking Differential Loss Algorithm
To prevent $1.25$ Billion float Vocabulary OOMs when applying $50\times$ differential penalties to label tokens, we utilize the following algorithm to iteratively compute loss within a bounded VRAM footprint.
Algorithm 1 OOM-Safe Micro-Chunking Weighted Loss
Input: Logits $\hat{Y}\in\mathbb{R}^{B\times S\times|V|}$, Labels $Y\in\mathbb{Z}^{B\times S}$, Weights $\mathcal{W}$, Chunk Size $c$
Output: $\mathcal{L}_{final}$ (Scalar)
1:Flatten shift-logits: $\hat{Y}_{flat}\leftarrow\hat{Y}[...,:-1,:].\text{view}(-1,|V|)$
2:Flatten shift-labels: $Y_{flat}\leftarrow Y[...,1:].\text{view}(-1)$
[…]
5:for $i=0$ to $T$ step $c$ do
6:  $\hat{y}_{chunk}\leftarrow\hat{Y}_{flat}[i:i+c]$
7:  $y_{chunk}\leftarrow Y_{flat}[i:i+c]$
8:  $Mask\leftarrow(y_{chunk}\neq-100)$
9:  if $\text{any}(Mask)$ then
10:   $L_{raw}\leftarrow\text{CrossEntropy}(\hat{y}_{chunk},y_{chunk},\text{reduce}=\text{none})$
11:   $w_{chunk}\leftarrow\mathcal{W}[y_{chunk}\cdot Mask]$ $\triangleright$ Safe weight lookup
12:   $L_{weighted}\leftarrow L_{raw}\times w_{chunk}$
13:   $L_{clamped}\leftarrow\min(L_{weighted},\>5.0\times\max(\mathcal{W}))$ $\triangleright$ Gradient Seesaw Fix
14:   $L_{total}\leftarrow L_{total}+\sum(L_{clamped}[Mask])$
15:   $W_{total}\leftarrow W_{total}+\sum(w_{chunk}[Mask])$
16:  end if
17:end for
18:return $L_{total}/\max(W_{total},1e^{-9})$
Appendix D VeNRA Prompting Physics (Bookending)
The Sentinel Reverse-CoT prompt relies on strict selective repetition to overcome the “Lost in the Middle” phenomenon over 4,0004{,}000 tokens.
<|im_start|>user
### TASK:
Verify if the CLAIMED_ANSWER given the AGENT_TRACE is [Found, Fake, General].
### VERIFICATION TARGET:
**Query:** {query}
**AGENT TRACE:** {trace}
**Claimed Answer:** {statement}
### EVIDENCE (Source Document):
{context_chunks}
### VERIFICATION TARGET Recap:
**Query:** {query}
**AGENT TRACE:** {trace}
**Claimed Answer:** {statement}
### AUDIT ALGORITHM (CRITICAL):
Do NOT recalculate the math. Assume the arithmetic evaluates to the Claimed
Answer. You must verify the EXTRACTION and LOGIC. Output your label first.
<|im_end|><|im_start|>assistant
Label: